OpenStudy (anonymous):

solve each system by substitution, show work Y = 6x - 11 -2x - 3y = -7

OpenStudy (anonymous):

so... -2x - 3 (6x - 11) = -7 and solve?

hero (hero):

Yes, solve for x.

hero (hero):

You still have to find y, but first solve for x.

OpenStudy (anonymous):

-2x - 18x - 11 = -7 ?

hero (hero):

No, you have to multiply -3(6x - 11) correctly

OpenStudy (anonymous):

erm ._.

OpenStudy (anonymous):

-2x - 18x - 33 = -7 :D

OpenStudy (anonymous):

my answer was x = 1.3